OpenStudy (anonymous):

integral (x^2 -x+2)/(x^3 -1) dx

OpenStudy (anonymous):

this a job for partial fractions ick

OpenStudy (anonymous):

yh

OpenStudy (anonymous):

somehow i am getting a positive with tan^-1 in the answer

OpenStudy (anonymous):

something like \[\frac{x^2-x+2}{x^3-1}=\frac{a}{x-1}+\frac{bx+c}{x^2+x+1}\] and then integrate that thing

OpenStudy (anonymous):

yeah i think \[\frac{bx+c}{x^2+x+1}\] will give an inverse tangent.

OpenStudy (anonymous):

yh i got inverse tan but with a positive(according to the book it is negative)